p-ADIC CHAOS AND RANDOM NUMBER GENERATION

CHRISTOPHER F. WOODCOCK AND NIGEL P. SMART

Abstract. We discuss the properties of p-adic analogues of the logistic and Smale horseshoe maps, and adapt them to form possible practical pseudo-random number generators. The properties of these practical modifications are studied in detail in the case p = 2.